OpenStudy (anonymous):

Find the mean, median, mode, mid-range, range, variance, standard deviation, coefficient of variation, and outliers (if any) of the following data set: 5, 10, 14, 16, 16, 21, 23, 33, 36, 44, 46, 53, 56, 62, 64, 66, 67, 77, 82, 83, 87, 88, 92, 93, 99, 100, 102, 105, 114, 120.

OpenStudy (anonymous):

More than anything I want to verify my numbers are correct. The results I came up with seemed kind of extreme. Thanks!

OpenStudy (e.mccormick):

Well, did you have any trouble with the standard formulas for these?

OpenStudy (texaschic101):

check yourself with this.. http://www.calculatorsoup.com/calculators/statistics/descriptivestatistics.php

OpenStudy (anonymous):

Not that I think of. Just extreme numbers so I could have missed something.

OpenStudy (e.mccormick):

The one issue with using random calculator sites is the two standard deviation formulas. Not all classes actually tell you if they are using the population or sample formula or what the difference is. \(\sqrt{\dfrac{\sum\limits_{i=1}^N(x_i-\overline{x})^2}{N}}\) vs \(\sqrt{\dfrac{\sum\limits_{i=1}^N(x_i-\overline{x})^2}{N-1}}\) So if the online calculator and the class use different ones, you can get the wrong answer. Though, what that site texaschic101 linked gives you should be close to the answers you got.

OpenStudy (anonymous):

Using that site, I pretty much got all the same results, with a slight difference in my Variance.

OpenStudy (e.mccormick):

Then you might just be using the other formula.

OpenStudy (anonymous):

It appears that they rounded their answer where I rounded to two decimal places. Still does not give any information on the possible outliers.